OpenStudy (anonymous):

Fan & Medal 4) Use the Polynomial Identity below to help you create a list of 10 Pythagorean Triples: (x^2 + y^2)^2 = (x^2 - y^2)^2 + (2xy)^2 Hint #1: c^2 = a^2 + b^2 Hint #2: pick 2 positive integers x and y, where x > y

OpenStudy (anonymous):

Would I be able to just list any 10 Pythagorean triples?

OpenStudy (anonymous):

I think you can pick x=2 and y=1, then x=3 and y=1, then x=3 and y=2 Things like that and it will give you the tripples

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

If you let x=2 and y=1, you get: \[ (2^2+1^2)^2 = (2^2-1^2)^2+(2(2)(1))^2 \implies 5^2=3^2+2^2 \]This is your first Pythagorean tipple.

OpenStudy (anonymous):

Ohhhh, that clears up a lot. Thanks. I'll start doing it and post what I get to see if it's right.

OpenStudy (anonymous):

Wait, I thought the lowest Pythagorean triple was 3,4,5?